A piston has an external pressure of 6.00 atm. How much work has been done in joules if the cylinder goes from a volume of 0.120 liters to 0.620 liters?
Answer #1

W = -Pext(Vf-Vi)
P = 6.0atm
Vi = 0.12 L
Vf = 0.62 L
put values in above expression
W = -Pext(Vf-Vi)
W = -6.0atm (0.62L - 0.12L)
W = -3.0 atm.L
W = -3.0 *101.33 J
W = -304 J
Answer:   -304 J